OpenStudy (anonymous):

I need help with my homework quiz. If f(x) = x^3 and g(x) = (x + 1)^3, which is the graph of g(x)? 1) a vertical transformation of f(x) 1 unit upward 2) a horizontal transformation of f(x) 1 unit to the left 3) a horizontal transformation of f(x) 1 unit to the right 4) a vertical transformation of f(x) 1 unit downward 5) a vertical transformation of f(x) 3 units downward

OpenStudy (anonymous):

What doyou think the answer is

OpenStudy (anonymous):

I believe the answer is #3 but im not sure, if someone could show me what to do for the problem ill do it.

OpenStudy (anonymous):

well the transformation is affecting the x variable and usually setting the x to zero will show you what to do with the coordinate so doing that what do you get

OpenStudy (anonymous):

well i was meaning x+1=0 so x=-1

OpenStudy (anonymous):

and you already established it was a horizontal transformation since it afects the x

OpenStudy (anonymous):

so transformation to the left since x=-1?

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

Thank you!(: Will you help me with my others?(:

OpenStudy (anonymous):

haha no problem ill do what i can xD

OpenStudy (anonymous):

If the function y = x^5 is transformed to y = x^5 + 3, which statement is true about the transformed function? 1) It is an even function. 2) It is an odd function. 3) It is neither an even nor an odd function. 4 )It is both an even and an odd function. 5) It is symmetric with respect to the x-axis.

OpenStudy (anonymous):

Replace x with -x and compare the result to f(x). If f(-x) = f(x), the function is even. If f(-x) = - f(x), the function is odd.

OpenStudy (anonymous):

so it would be even?

OpenStudy (anonymous):

woops wrong q

OpenStudy (anonymous):

well it wouldnt be even because solving with a -x does not give the same result

OpenStudy (anonymous):

the function is exactly the same except for a 3 added to the y which would be a vertical transformation

OpenStudy (anonymous):

so i think it would be 3

OpenStudy (anonymous):

ohh since it adds 3 and thats for the Y and not the X.

OpenStudy (anonymous):

yes